250 WHT – Booster Pump Replacement

01/22/20 – TEI fieldwork visit and researching equipment options.
01/30/20 – SOSS submitted and original fieldwork completed.
02/24/20 – SOSS submitted awaiting approvals.
03/23/20 – Project ready for review. TTC will send to county for review. New booster pump is 100% redundant but new pumps are 5HP. Barth confirmed conduit and feeders still sized properly for new loads.
04/16/20 – Bids received from Peck, Geiler, Triton and Debra. TEI tabulated and sent to the County for decision.
04/28/20 – TTC released bid tab to bidders. Did not tell bidders who the winning bidder is.
04/30/20 – PO in for approval with County, looking at weekend work for this job.
05/27/20 – Project in for approval with the County for weekend work with PHB as the apparent low bidder.
06/24/20 – PHB sent the submittals in on the project. ThermalTech reviewed and sent back to the vendor including limited comments from the County. PHB working on background checks with BC who is helping get them onsite to the work ASAP. Unfortunately, the lead time on the pump package is 5 weeks at this time because we missed the “on shelf” package they had waiting for us. The hope is to get this installed ASAP.
07/24/20 – PHB has scheduled the pump replacement on August 8 with the backup date of August 9. They have not however received their PO and Contract yet. TEI check with Facilities and it is in place and Tony will send it. PHB will install the pumps and get them working on the weekend BAS interface and local alarms will be picked up the next week by Peck.
08/26/20 – Peck performed the work over the 8/8 weekend. The new booster pump is working and online. TEI issued Punchlist for the project to PHB with remaining work items. Payapp #1 was submitted and approved. Final payapp was submitted but held for closeout documents.
09/16/20 – TTE on-site performing Records Drawing check. Some piping needs to be changed on Record Drawing and CAFM drawing.
09/18/20 – OM’s delivered. Punchlist work complete. TTE to approve the final invoice and send in the final paperwork. Project ready to archive.
10/28/20 – TTC checked in with PHB to ensure all payments have been made and project can be closed.

by Champlin Champlin

250 WHT – First Floor Painting

10/16/19 – County made request for get pricing to bid out painting 250 William Howard Taft – first floor only.
10/25/19 – Champlin was at site for field verification.
12/5/19 – Cost Estimate and SD drawings sent to County
12/10/19 – Cost Estimate Revised
01/9/20 – SOSS submitted to county
01/22/20 – Approved SOSS and PO received by Champlin
01/30/20 – Waiting to receive bid schedule approval and list of approved vendors before sending RFP to potential bidders
02/21/20 – bid documents sent to County for approval
03/11/21 – Documents revisions sent to Champlin
03/18/20 – Bid documents sent to county
04/02/20 – Project put on hold due to Covid-19
06/25/20 – County reported that most of this work was completed in house.

by Champlin Champlin

250 WHT – Egress Door

06/18/19 – Bert emailed regrading a question from the County’s Safety Specialist regarding a egress door and egress path at the 250 WHT Building.
06/18/19 – Bert emailed floorplan and egress door in question to Champlin for review prior to site visit.
06/25/19 – On site meeting with Bert, Tony Matre and Kru to review location of Exit Door in question and review egress path.
07/03/19 – John from THP shared building code notes from 2015 project for replacement of deteriorated concrete bridges.
09/01/19 – Champlin shared memo explaining removing exit door designation at mid-landing.
09/11/19 – Cincinnati Fire Department sent followup questions.
09/17/19 – Champlin reviewed existing hardware and access control with building manager. Champlin to review this with Fire Department.

10/24/19 – Champlin had a code review meeting with Cincinnati Fire Department (Darryl Meadows) and Building Department (Greg Wiles)


250 WHT – Trench Drain Investigation (F115)

09/20/18 – County requested a SOSS to investigate the south trench drain and provide recommendations to repair.
03/21/19 – THP visited the site. THP to review conditions with County.
03/21/19 – THP visited the site. THP to review conditions with County.
04/05/19 – THP completed a follow up to visit the Trench drain.
04/25/19 – Follow meeting with Bert work on hold.
09/26/19 – Project ready to archive.


250 WHT – Access Control System Code Review

02/20/18 – Thermaltech meets with Mike Jackson and Tony Matre to discuss needs for code research for the existing access control system at building.

03/27/18 – ThermalTech to send Tony Matre a draft copy of the code research report by end of March for review.

04/25/18 – ThermalTech sent the report to Tony Matre.   Study complete.


250 WHT – Asbestos Study

Project Lead: GBBN

10/20/11 – Kick-off Meeting with Pinnacle and Facilities Department

11/04/11 – Facilities Department sens list of buildings to be included in Asbestos Abatement Reporting.

02/06/12 – Pinnacle Engineering requested drawing files for the purpose of documenting findings.

02/07/12 – GBBN sent drawing files to Pinnacle Engineering in PDF form.

03/30/12 – Meeting with Tim Schue of Facilities Department to discuss Archibus Input Strategy

04/05/12 – Meeting with GBBN and Facilities to discuss drawing format.

05/10/12 – Meeting with GBBN and Pinnacle to discuss drawing format for reports. Tony Matre’s request for ‘General Information Notes to occur on each floor plan are discussed.

01/26/13 – BRG confirms entry of ACM data into Archibus is completed.

01/28/13 – GBBN conducts spot review and confirms data entry is accurate.

04/05/13 – Email sent to Pinnacle Engineering inquiring about status of project reports and unmet schedule milestone goes without response.

04/20/13 – Email sent to Pinnacle Engineering inquiring about status of project reports and unmet schedule milestone goes without response. Reported to Tony Matre via telephone call.

02/24/14 – GBBN and Pinnacle Engineering met with Tony Matre to deliver the final reports and electronic copies of the reports. GBBN is completing final report drawings for Pinnacle as final preliminary drawings are delivered by Pinnacle to GBBN. All drawings are to be completed by 2/28/14.

03/26/14 – All asbestos reports and associated drawings are complete. Final drawing input into Archibus/WebCentral is ongoing. GBBN was granted access to the database to complete the drawing downloads.

07/01/14 – GBBN sends another email and follow-up voice mail to Pinnacle Engineering inquiring about status of billings and final completion.

8/29/14- Pinnacle Engineering sends invoices for services to GBBN.

10/29/14 – GBBN sends Tony Matre SOSS invoice status with all of the Pinnacle Engineering invoices included.

10/31/14 – GBBN pursuant to Tony Matre’s request sends letter of documentation regarding scope overruns.

02/20/15 – GBBN in communications to finalize the remaining invoices to close out these projects.


250 WHT – Plumbing Division Upgrade

06/14/05 – Phasor turned in a new change order for Fire Alarm additional work at 250. The Change Order cost including all the additional inspections is $967.82. TE will approve and send to Diana for signing. Please return a signed copy to both ThermalTech and Phasor when complete.

06/08/05 – TTC sent final Reupert paperwork to Diana for complete project closeout. Noise issued on floor is being address by Mark Hardin.

05/11/05 – Sound issues arose from the toilet room and accompanying space being too quiet. GBBN made a field visit to investigate and the consultants gave recommendations about white noise or another exhaust fan. The County is to advise on how to proceed from here.

05/10/05 – TE emailed GBBN to advise that toilet exhaust on this project is connected to the main building toilet exhaust. If there is a problem have Tom check the main fan and if that doesn’t work, TE will measure the airflow and check the existing conditions contingent on a future request by the County.

05/02/05 – The building passed it fire alarm inspection this morning at 6:00 a.m. TTC, Jack Lamping, Darnell, Jim McFarland and Darryl Meadows were present and witnessed the test and the paperwork was signed. The City Inspector had issues with the existing system but did not hold up or not pass the test but these are issues, he wishes the County would fix because the system is technically not installed properly. Item (1) any smoke or pull device on the first floor reports as “Smoke & Pull Stations”. The inspector strongly urges that these two actuation devices be separated and report separately when in alarm. Item (2) when in alarm and the “Alarm Silence” button is pushed it disengage all the horns AND strobes on the first floor – this is against code as ONLY the horns should be disabled and the strobes should continue to flash until the system is RESET (it was shown many times during our testing periods that they are not wired properly). Item (3) They questioned the “fail-secure” status of the doors upstairs and downstairs. When the FA system is in alarm all the door holds release as they should and all the security access is disabled so that a key is required to enter the space. I confirmed with TA and MS that this is the correct operation for your building. The people inside can always egress out as there are strike bars or non locking door handles from the inside to the outside so none of that is required to change. Item (4) The first floor system is a “System Sentry” type system which was in place when Simplex received the second contract to renovate the second floor space. They did this by daisy chaining all the existing 4-wire devices through a ZAM (Zone Actuation Module) which is why they all report as the same device back at the Simplex panel. TE would recommend putting in a project right now that upgrades that existing system to 100% Simplex (which is a fully addressable system 4010 model). We estimate about a $23,700 project (30 devices @ $9000, 3000LF @ $9000, 6 ZAM’s @ $3000, Programming @ $1500, Permit and Testing @ $1200). Please advise.

04/28/05 -TE spent most of the day at the building trying to get the Simplex system working in conjunction with the System Sensor system. We tried many custom programs but in the end nothing worked perfectly. TTC has 3 ways to try to get system fixed this morning and will meet Simplex back on site to repair the system. (1) write a custom program that actually works, (2) replace the 4 smoke detectors with authentic Simplex devices, (3) add a UL listed manual reset to reset the 1st and Bsmt floors. TTC will keep DES up to date as we go through the day.

04/27/05 – The project failed the initial Fire Alarm test this morning. Not from device malfunction but the actual FA panel would not reset after receiving an alarm from a smoke head. The system was pre-tested using pull stations and reset fine on Tuesday. Because of this the City did not issue an approval. Phasor, TE and the County contacted Simplex for an emergency call into a technicians to repair or reprogram the main FA panel. Jack talked to Simplex and will be back on site as soon as they arrive today to work on the system to assist in figuring out what is going on and wrong. He believes that he was able to get the technician to understand what was happening and hopes to resolve the issue today. The inspectors agreed to come back this week to approve.

04/26/05 – Phasor bought two strobes locally and will install them today and have everything ready for tomorrow morning. Jack said it was a scramble but he found them and is ready to install. They will be ready for tomorrow’s test. At 2:48 p.m. they reported that everything is set for final FA inspection on Wed at 6:00 a.m.

04/19/05 – Final fire alarm test is scheduled for 04/27/05 at 6:00 a.m. Tom will be on hand from the County for FA issues. TE will attend as Owner’s Rep. Phasor will be EC. Copp stated that all egress issues are resolved and will be working fine for our final test.

04/15/05 – TE contacted Phasor about final completion and inspection by email to make sure everyone is ready for next week.

04/06/05 – TE and Phasor and County are coordinating final inspection details for April 25th. TE waiting on contract interior information from County before confirming

04/05/05 – Project Meeting today. Set new schedule for completion. JB Schmidt scheduled to be complete by April 15. Phasor to be work on 4/15 and 4/18-4/20. Final Inspections to be tentatively scheduled for 04/25/05. TE to contact Phasor for FA inspection coordination with JB. TE to contact Copp about schedule for plumbing space and general building space and to work on resolving software issues. County ES/Facilities to contact Sheriff’s Dept about Panic button for receptionist in ES.

03/25/05 – Reupert delivered some of the final paperwork with no payapp today. TE will contact them about this.

03/15/05 – Portions of the project have been on hold awaiting the replacement of drywall and water sealant. TE received final paperwork and O&M manuals including balance reports from Reupert. The manuals are not bound or copied. TE will copy and make sets for County use and turn them in at the end of the project with final paperwork and payapp from HVAC contractor.

01/31/05 – TE attended last Thursday’s Const Mtg. HVAC is complete. EC is nearly complete. PC is complete with working fixtures. Hot water reached the LAV in under 60 seconds. All fixtures appeared to drain properly. Security is nearly complete but Copp has some final work to do.

01/26/05 – TE faxed one copy and mailed three copies of the Affidavit of Release of Liens to Reupert.

01/24/05 – Jamie of Reupert called TE and indicated that she was preparing final document packet but did not know where to find the Affidavit of Release of Liens forms. TE to mail the Affidavit of Release of Liens to Reupert.

01/18/05 – TTC sent Reupert their punchlist item (balance report) and asked for their O&M’s.

12/27/04 – TTC stopped by the building to coordinate lighting layout with Jack Lamping. Some of the HVAC piping is too close to the ceiling grid and the deep parabolics don’t fit into the grid and below the ductwork and 3-way heating piping and valves. We remedied this situation by moving some lights over a grid space or 2. TTC has a copy of the new locations for as-built purposes. TTC believes the new layout will provide sufficient lighting levels and be a workable layout for the building occupants. The only outstanding issue with the lighting at this point is that the light nearest the column near the restroom can either be swapped out with a shallower fixture like the ones slated for the storage room and the lens be replaced with a paracube or (2) 2×2 fixtures can be used instead of (1) 2×4 fixture. The former would cost the project little or no money, the ladder may cost $200 – $300 out of the remaining allowance. We should have an a final answer later today.

12/20/04 – Ken Reupert out of the office, Jamie will have both of the Affidavit of Original or Subcontractors sent to ThermalTech today or tomorrow (12/21/04).

12/09/04 – Reupert revised Pay Applications 1 & 2 received in the mail. Voice message left for Ken Reupert to disregard earlier message and that the Pay Application will be held for the “Affidavit of Original or Subcontractor” form to be submitted.

12/09/04 – Reupert Pay Applications 1 & 2 Voided due to $600.00 allowance shown on application as a Change Order. Voice message left for Ken Reupert explaining that the Applications were incorrect and will need to be revised.

12/03/04 – TTC talked to Reupert today. All the above ceiling ductwork is complete at this point. The 3-way valves are all piped in. The controls on on schedule at this point. The GC will contact the MC when the diffusers and final connections are ready. The MC has to attach some angles around the fire dampers to meet code and should take of that work next week.

11/22/04 – Reupert called today to discuss the 3-way valves ordered on this project via the submittal change. TE instructed them to pipe the valves in as 3-way valves and use up to $1250 of the allowance for this work and to make the necessary scheduling changes to make this happen. Per our submittals to the MC and CC we changed the 2-way valves to 3-way because this building has constant speed pumps and to prevent dead heading the heating hot water system will require 3-way valves in most installs.

11/18/04 – Reupert called to ask for a size change in the return grill. TE approved a 24 x 12 return opening above the ceiling for this project. TE to verify on site at today’s meeting all their request

10/27/04 – Reupert submitted a $600 allowance request to rework the piping at the bulkhead pursuant to our last construction meeting. TE approved the request from the $2500 mechanical allowance for this project, new balance $1900.

Change order request in approved for security change order and is the amount is not being subtracted from the allowance. TEI issued a revised sketch that corrected the access controlled door locations.

10/21/04 – TE contacted Reupert about missing today’s construction mtg, making the next meeting for sure, submitting their schedule of values, change order pricing for relocating piping and VAV box submittals. Everything is in the works. Reupert is going to send us the schedule, the VAV boxes are already in their shop (we now have to approve them (don’t worry they will be fine)), they will begin working on the change order request.

10/13/04 – AIA Change Order for $1,854.46 mailed to Phasor

10/11/04 – The Plumbing Permit Number is: # 2004P07058.

10/08/04 – TEI to prepare electrical bulletin #1 for 1854.36

10/07/04 – GBBN provided the permit number for the project: The permit number is as follows: #2004P05650

10/05/04 – TEI received devices submittal via email from Phasor.

10/05/04 – DES confirmed that TE sent Lisa the stamped drawings for permit. TE sent them on Thursday 10/07/04.

09/29/04 – Jeanette provided a dimensioned drawing for HamCo plumbing to use for this project – TE reviewed and approved for use, TE to pickup plumbing fixtures from Bo for approval, TE to provide stamped drawings for permit use to Lisa, TE to coordinate revised security (Copp to upgrade entire building including 2 new doors) and TE to write change order for this work to Phasor, Reupert was not at meeting and TTC to confirm that Reupert will be at next meeting, TE requested that GBBN reschedule Oct 20 mtg as it conflicts with CH Archive mtg.

09/20/04 – Ben has asked TTC to coordinate water closet locations with Bo Blank.

09/14/04 – Copp is providing a price to upgrade the whole 250 building, this has to be coordinated through TE and Phasor since our electrical contractor has security money in their bid to do this work.

09/08/04 – Bo advised to remove sanitary so that Reupert can begin construction.

09/08/04 – Pre Construction Mtg completed.

09/03/04 – DES advised that contracts are approved. GBBN will setup pre-construction meeting.

09/28/04 – Lighting submittal approved 9/22/04

10/01/04 – Request for change order pricing from Reupert for Fire Damper

10/02/04 – Reupert gave a verbal price of $375 for the addition of a fire damper


250 WHT – Electric Device Study

10/18/06 – Final close-out meeting was held on location. Queen City has completed all contracted work. I have the drawings and will have the information put into electronic format over the next few weeks. I also took the panel index binders. I will have all documents delivered at the same time.

09/19/06 – Construction meeting was held on location. All field work was to be completed by this afternoon. Dick Kuhn of Queen City will be in contact with me when he if finished with his circuit verification. When I have been contacted I will contact all necessary persons to set up the close-out meeting.

08/28/06 – Kick-off meeting held on location.

07/18/06 – Contracts have been submitted.

06/19/06 – Pre-bid meeting was held on location.


250 WHT – Security Camera Upgrade

7/31/08 – Bert reported that the project is finished and final paperwork has been received. Final payment is approved to the contractor. Broken camera has been fixed. Contractor is awaiting final payment.
07/30/08 – Project is complete. County has all final paperwork.
05/16/08 – Bert sent back changer order and payapp #3 signed by County.
05/02/08 – Stapleton has submitted all final paperwork and sent back a signed change order on this project. TEI has reviewed and approved all closeout documents on this project. They are being UPS to Bert Watts for closeout on this project. The change order has to be signed by the owner and copies sent back to TEI and Stapleton for our records.
03/06/08 – Stapleton and Copp have reached substantial completion. The translator has been installed. Remaining items on the job include: our punchlist (early next week), training for the county, programming at the sheriff’s office, Copp is going to replace one camera that is acting up.
02/08/08 – Stapleton sent the change order cost to TEI for the translator and TEI passed it along to Bert Watts for approval on this project. TEI updated the website with current CO and Allowance costs.
01/23/08 – TTC requested a const meeting for this project this Thursday or Friday.
01/22/08 – MAH met with Dick (Stapleton) and Paul (Copp) about the mounting of the exterior cameras. Copp will supply (3) corner mount brackets for $267 allowance. Stapleton will install camera mounts when they arrive (should arrive 1/23). Also discussed was the data port being included on same plate as existing phone jack in lobby, lobby camera location (NE corner and viewing doors), and location of exterior camera at rear entrance door (will be mounted on adjacent wall approximately 24″ above window to avoid washout from wallpack light fixture).
12/14/07 – Stapleton discovered that the holes in the pole bases are only 2-1/2″ and will not accommodate (3) 1″ conduits as the design calls for. Stapleton requests a $500 change order to install junction boxes in the pole bases to solve this problem. TEI agrees with this solution and the change order amount.
12/12/07 – MAH clarified for Stapleton that the control signal for the PTZ cameras was sent over the coax video cable so no additional control wiring is necessary. Also, siamese RG-59 wire with 18AWG power is acceptable for all interior and building mounted cameras. A separate coax and #12AWG power pair is required for the pole mounted cameras that is wet/direct sunlight rated. Copp Systems submittals are approved and should be sent back 12/13/07.
12/04/07 – Const Mtg #2 was held today and TEI distributed meeting minutes. Stapleton will be staring site work this Friday. Ken Edgell will advise the tenants of the 250 WHT building. Stapleton will be boring the light pole bases and conduit routes this weekend. TEI, Ken and Stapleton pick light pole locations in the field for this work.
11/29/07 – The pre construction meeting was held last week. GM typed up the meeting notes. Stapleton called with questions about submittals, dwgs and const mtg time. TEI setup another const mtg for next week with the owner and facilities. We will discuss the RFI’s about the lamp bases and base diameter. Bert sent out the list of approved employees for this project today which included Barnett, Ronald J., Eckart, Carl, Hammitt, Wayne, Leopald, Mike, Leopold, Douglas A., Stevens, Brian K. Terry Cannon emailed Stapleton to ask about Copp system background checks. Mike Huntwork approved all submittals with No Exceptions noted.
10/03/07 – TTC sent the contract for Stapleton to Bert today via City Dash with an allowance of $2500 added to the project. If acceptable, the County can begin the contract process. Stapleton won the bid with a bid of $90,586.00 + $2500 in allowance money. Our electrical plan review was approved on 09/11/2207 with an application # of CIC0703025. The contractor will now have to pay and pick this permit up when the job begins.
09/10/07 – TEI is submitting plans to IBI for plan review and permit # for this job. TEI will provide permit numbers to County for record keeping purposes when we are complete with this process. It will be the contractors responsibility to pick up and pay for the balance of the permit when one is assigned for this project.
08/27/07 – This project has been sent out for bid with hard copies to Mark Donnelly.
08/24/07 – TEI has been instructed to move forward with this project. This project is set to open for bid on 09/26/07. The pre-bid meeting has been set up for 09/11/07.
08/10/07 – A project review meeting was held at the 250 WHT building on Tuesday, August 7, 2007. Items discussed included camera locations and specifications, new light pole locations, access into building, and budget. TEI is awaiting word from the Hamilton County Facilities Department on when this project will go out to bid.
07/30/07 – The drawings and specs are completed. The package is to be sent out today. There will not be front end or ITB# for this project because construction has been postponed until next fiscal year. When TE is given the go ahead from the Hamilton County Facilities Department on this project they will put together the front end of the project manual and issue for bid at that time.
07/09/07 – TE is finalizing the design. Drawings will be turned over to Hamilton County Facilities Department when complete.
06/29/07 – At the monthly Partnering Meeting with Hamilton County it was decided that the money to perform this work would be budgeted for a 2008 project. TE was directed to complete the design and turn over the drawings to the Hamilton County Facilities Department.
06/22/07 – A new cost estimate was presented to Mark Donnelly. TE is awaiting response on whether or not to proceed with the new latest layout and additional cost.
06/15/07 – Project moving along as expected. Changes per meeting with Mark Donnelly are being incorporated and a new cost estimated is being generated.
06/08/07 – New parking lot lighting and camera layouts were presented to Mark Donnelly along with a cost estimate. TE is moving forward with this project. There are a few slight adjustments; number of new poles, location of new poles, and number of cameras. These modifications will be made and presented again to Mark Donnelly.
06/01/07 – TE will present Mark Donnelly with new parking lot lighting and camera layouts early next week. A cost estimate will be included with the new layouts.
05/22/07 – Project review meeting was held today. Various points were discussed as described in the meeting notes. TE is laying out different schemes of possible lighting and camera layouts with the points discussed in the meeting kept in mind.
05/21/07 – Meeting to review current status is schedule for May 22, 2007 with Mark Donnelly.
04/27/07 – Estimates are complete and will be forwarded to Hamilton County Facilities Department today, April 27, 2007.
04/20/07 – Estimates are near completion and should be turned over the Hamilton County Facilities Department by the end of next week, April 27 2007.
04/13/07 – TE is working on the estimates to be turned over to Hamilton County Facilities Department.
04/06/07 – Review meeting was held today, April 6, 2007. TE is to provide pricing on 2 systems and then await owner’s selection.
03/30/07 – Preliminary layouts have been presented to Hamilton County Facilities Department for review. Review meeting will be scheduled for the first week of April 2007.
03/20/07 – TE met with camera vendor. Parking lot lighting and camera layout is currently underway.
03/20/07 – Design is continuing. Parking lot lighting is under design.
03/14/07 – NR created project journal and filled in back data from previous meetings.02/14/07 – Proposed parking lot lighting and camera locations has been completed by TE. This information is awaiting review by Hamilton County Facilities Department.
02/28/07 – Foot candle readings were taken by Phil Lambing in the evening of this date.
02/27/07 – Kick-off meeting held on location by Phil Lambing of ThermalTech. Attendees include Mark Hardin and Bill Scholl of Hamilton County Facilities Department. IT was decided that TE will collect foot candle readings in the parking lot to get a good idea of what will be needed to ensure personnel safety


250 WHT – Monitoring Room Upgrade

Current Status: Construction

03/30/11 – This project is complete. TEI finalized the HVAC permit and paid for it at the City of Cincinnati so EDI can close out this project.

03/02/11 – TTC checked on braided hose install by EDI on humidifier. Brady said they are working on this. TTC review wall mounted installation and there is no good way to clean this install up with a simple product. TTC asked EDI to closeout project and send in final payapp.

12/30/11 – TEI released the punchlist on the project today.

12/15/10 – EDI on board and working on project towards punchlist completion


250 WHT – Phone Switch Room

Current Status: Closeout

Contract Date: xx/xx/11

Closeout Status: Requested

06/07/11 – Kick-off Meeting at 250 WHT Building. The existing phone switch room at 237 WHT was also visited.

06/15/11 – SOSS delivered to Bert Watts.

06/23/11 – Signed SOSS received with associated Purchase Order.

06/28/11 – Design has begun. There are concerns about the fire suppression on this project already. Wet sprinkler exist in the storage rooms and IT wants to remove them. TEI says that the city would have to approval such a removal. There are discussions about Alternate cost for a clean agent fire suppression system for this space but TEI is also recommending an alternate to that to provide a pre-action system if we have to keep the wet sprinklers.

07/06/11 – Kick-off Meeting with Mike Barth and Mike Folck of ThermalTech. ThermalTech and GBBN held coordination meeting.

07/11/11 – TTC has suggested asking the city if we can remove the spinklers that existing in the storage rooms now that are limited area sprinklers.

07/21/11 – ThermalTech, HamCo facilities, building personnel, and HamCo IT staff met to review progress drawings. HamCo IT group modified plans to change the location of phone racks, phone switches, voice mail server, and associated equipment. The location of the condensing unit was determined and the routing of the line set was planned thru the existing elevator shaft. Several alternates were discussed including a clean agent fire suppression system and removal of the existing freight elevator.

07/29/11 – GBBN found a transfer grille that ThermalTech will have to address in the clean agent alternate before bidding.

08/05/11 – HamCo facilities and HamCo IT review comments incorporated into design. Drawings and specs issued for bidding including clean agent fire suppression and elevator removal alternates.

08/08 /11 -Bid documents released to Purchasing and ARC Reproductions.

08/18 /11 -Pre-bid meeting held at 250 WHT. Deadline for questions: 08/22/11, deadline for answers: 08/25/11, bid opening date 09/01/11.

08/24/11 – Addenda #1 released by GBBN to Purchasing and ARC Reproductions. Issued to bidders for clarification.

09/01/11 – Bid Opening held. Contractors are selected.

09/02/11 – GBBN authors contract for General Trades Contractor (JB Schmitt Co., Inc.)

09/05/11 – TEI suggests revisions to contracts. Bert Watts approves those revisions to be made. GBBN sends modified contract and signature sheet to Bert Watts.03/27/12 – ThermalTech added new panel PP-1 to the Arc Flash model. An Arc Flash label was added to the panel in the field, and a copy of the updated excel file was emailed to Rodney Lofland and Bert Watts.

10/19/11 – Received Contracts and Purchase Orders from Leo J. Brielmaier Co., J.B.Schmitt Co. and Atkins and Stang Inc.

12/05/11 – Pre-construction meeting held with TEI, GBBN, HamCo, and contractors to review status and coordination. GC and EC expected to mobilize tomorrow (12/6), MC to mobilize when completed with background checks.

12/13/11 – Concrete pad size and location within telephone switch room approved by TEI and GBBN. J.B. Schmitt to proceed with installation.

02/22/12 – Project is substantially complete, waiting on close out documentation. There is a need to have a conduit extended from the existing telephone room in the 237 building to the TB clinic. TE to hold on any work associated with this until further notice from Bert Watts. Upon receipt of all closeout documentation from all vendors, TE will forward to County.

02/23/12 – Bert reported that telecom will be changing the way the fiber is running but County telcomm will coordination this work directly with Neyer and Cinbell

02/28/12 – ThermalTech is in receipt of the closeout documents that you were submitted for this project, along with a receipt for final payment. Per the requirements of the project manual, there are items that must be submitted prior to processing the final payments. Upon receipt of the following items, we can process the closeout documents: Punchlist (completed and signed), Notarized statements of warrantee, AIA Document G706 Contractors Affidavit of payment of debts and claims, AIA Document G706A Contractor’s affidavit of release of leans, separate affidavits and waiver of liens for each subcontractor and material supplier, AIA Document G707-consent of surety to final payment, affidavit of Contractor prevailing wages for all contractors or subcontractors supplying labor to the project.

4/4/12- GBBN As-buiilt drawings were sent to Bert Watts.

05/29/12 – As-builts were sent to Bert Watts.

06/26/12 – Mike Barth of ThermalTech contacted Fred Stang of Adkins and Stang to request all final closeout documentation. Fred indicated that Linda McCormick of A&S on vacation this week, but will follow up ASAP.

07/25/12 – Mike Barth spoke to Linda McCormick of A&S. She indicated that she has the entire close out package ready to submit except the as-built. Linda told me that she would get that to submit as soon as possible, and that they realize that the final pay app is still pending.

8/29/12 – Fred Stang of Adkins and Stang emailed on 8/14/12 indicating that closeout documents would be sent by the end of the week of 8/13/12. To date, still pending.


250 WHT – Boiler Replacement Project

Current Status: Closeout

Contract Date:

Closeout: Feldkamp has submitted all closeout paperwork. Glenwood has submitted all closeout paperwork.

08/13/10 – We are waiting on an approved SOSS to proceed.

09/24/10 – Project was bid out as scheduled. Bid opening scheduled for Oct 14.

10/18/10 – ThermalTech has recommended Feldkamp and Glenwood. Contracts are waiting on Pre Award responses from Feldkamp. There are questions about the controls contractor.

01/24/11 – Contracts are signed. Tony said he is ready to have the pre const mtg.

02/22/10 – Pre Const mtg complete. Electrical submittals have been sent in and approved. Pre Const Mtg notes sent being completed and sent to Contractors. We are waiting on submittals for the boilers. New project schedule must be approved by Owner.

03/30/11 – TEI has sent submittals back on all the mechanical and all the electrical submittals received. We are still waiting on flue piping submittal and controls submittals. Boiler equipment is about 4 to 6 weeks away from delivery.

04/27/11 – Asbestos abatement going on right now on some equipment. Const set to begin June 1. We are waiting for a schedule from Feldkamp and submittals on controls.

05/03/11 – Tony replied with the following “I would add one more item, that would be background Checks. I have received the (4) Local Checks. The remaining personnel on the list of employees that will be working at this site will need to have a background checks? I have not received any more Local Checks or any BCI Checks.
No work can be started until all the background are complete.

05/03/11 – TEI contacted Feldkamp and asked what is the status of this project? We don’t have project schedule, boiler flue submittals, and control submittals? TTC asked Tony to contact Feldkamp about project status.

05/04/11 – Terry wrote to the Team that Feldkamp had submitted their project schedule and it looks like this: June 1-Project start up, June 1-Demo, June 1-Electrical Demo & Work, June 14-15- Set Boilers, June 16-24-Pipe Boilers, June 16-24-Conrols, June 27-Flue piping, June 29-Inspections, June 30-Factory Startup. Basically the project construction side will take one month. To be safe, I would plan on another month to work out construction and control issues and plan on having something ready on July 31 or so.

05/24/11 – Control submittals returned. Const ready to begin in June but I don’t think contractors will have all background checks complete.

06/22/11 – TTC asked Pat Heeney Feldkamp to provide the pump vendor with pump head or whatever he thinks is best and get them the order.

06/28/11 – Feldkamp has made good progress on the demo since starting. We asked for a change order on the chilled water pumps. They said that they would be back this week to cleanup the boiler room and start to get ready for new equipment.

07/11/11 – ThermalTech checked out boiler progress and change order work on the chilled water pumps for this project. Boilers are demo’ed fine. 2 VFD’s have been installed for the HHW pumps, OA has been rerouted. Feldkamp has fitter change and we are trying to setup a const meeting for the week of July 26.

07/21/11 – ThermalTech was on site to review status. New boilers have been set in place, but no connections have been made. No indication of any contractors on site.

07/29/11 – Const Mtg 4 was held last Tuesday. The schedule was updated to have the work completed by Sept 2, 2011. Feldkamp has a new pipe fitter for this job. We discussed change order opportunities for the chilled water pumps. Vendors are giving us prices. Feldkamp has a $5000 allowance, Glenwood has a $2500 we can use. The exhaust vents will be routed straight up from the area well with rain velocity caps.

08/12/11 – TTC spoke with foreman regarding piping layout, discussion is as follows:

1. The old boilers were individually connected to the make-up water supply with no backflow preventer and we did not catch the fact that the piping only had a regulator inline before the boilers. You will need to provide a Watts 009 type domestic water backflow preventer to meet OPC section 608.3 to meet Code. Please do the work now and give me a price towards the allowance that I need to use.

2. On the pumps the P&ID line sizes here are the updates. The main line size is shown as 6″ but really most of piping at the pump is 4″. I will update the as-built.

3. On the inlet side of the pump to the suction diffuser that is coming down as 4″ piping and will not require a reduce as we show on the dwg as the suction diffuser can bolt right to the pump inlet at 3”.

4. On the outlet side of the pump the flex connectors came in at 4″ but we don’t have 4″ on the outlet. One pump has 3″ pipe and one pump as 21/2″ pipe. The flex connectors will be size to match the existing piping in the field.

5. The triple duty valves that were existing on the old pumps are not shown to be reinstalled on the new pump. This is correct, we are not requiring the triple duty valves because these pumps are being retrofitted with VFD’s. What we are requiring is what is shown on the PID’s as new butterfly isolation with lockout capability strainer and reducer if needed but the suction diffuser takes care of that. The flex connections are on the outlets and get sized to match.

6. Finally HWP-1 and HWP-2 are the same size pump in the new scheme but one has 3″ outlet piping and one has 2.5″, Thermaltech will update the as-built to reflect this change.

7. Feldkamp showed me how they are going to hook up the MER AHU unit back to the heating supply and that looked like an acceptable plan as well.

The gas piping looks great. The HHW piping at the boiers looks great.

08/22/11 – The Power terminations to the equipment at the 250 boiler job is complete. We still have to perform a factory authorized startup on the variable frequency drives. The control wiring does not have to be complete for us to do the startup but there should be a control person present. We require 1 week notice of when we can perform this startup.

08/25/11 -A&M need access to do controls work. Question into Tony about them starting now with the local backgrounds complete.

08/30/11 – Feldkamp provided a price of $45,000 to replace both CHWP-1 and CWP-1 today via a quote from his vendor. We only $5000 in allowance money so I advised Pat that it would be up to Tony Matre but I did not think he would accept this price for these pumps. I suggest we get quotes from other vendors if needed. Feldkamp says state inspector is complete, flow switches need to be added before he will sign off on install. Controls and startup should be complete by 9/8 or 9/9.

09/14/11 – Modbus is up and running per specs, and A&M has installed code to covert form Centrigrade to Fahrenheit. Boiler and Pump controls programming is complete, A&M will be testing it today. DDC Controls are about 95% complete at this time. Duke is supposed to contact TTC about regulator visit to site. Contractors have punchlist items to complete in 14 days from now.

09/15/11 – Duke is going to meter the incoming gas pressure before and after the meter and report pressure trends to TEI. The building is supplied by what is called “standard” gas service which meant 4-7″ pressure most of the time. It can drop lower in cold weather. When the data is received from Duke a decision will be made about what to do. Duke was also asked to check the regulator inside the building and was told his records indicated there was no regulator at this site-which means the regulator that is high on the wall could be an Owner one and not Duke, which means it could be removed and put boiler regulators back on. This could probably alleviate the pressure problem if the house regulator is what is causing the issue. The only problem is to check the pressure for the water heater that the gas line feeds, or to adjust the house regulator to make sure it gets the least pressure drop through the unit as possible. Once Duke is on site, a final decision can be made. Duke will be out Friday or Monday to install the trending device.

09/22/11 – All closeout paperwork has been received by ThermalTech. TTC processed it and is planning on dropping it off to Tony Matre today at our 2pm LEED meeting with GBBN.

09/23/11 – Close out Documents have been received by both Contractors.

09/30/11 – Tony is paying out Glenwood in full minus allowances but holding Feldkamp closeout for boiler resolution.

10/13/11 – Gas pressure and flow should be fine. The Duke readings were 7.9″ at full flow. Several issues to note on this job are:

1. Our boilers are smaller and therefore there is no real way that Duke can say our gas mains are undersized.

2. There is very little gas piping in this building therefore the overall gas pressure drop is very small even if Duke thinks the 1″ lines to the boilers should be 1,5″.

3. Boiler 1 is running and continues to run just fine.

4. The problem is in the BLR-2 machine itself and we should pursue Turton and tell him the gas pressure read by Duke is 7″ and ask him is there any way to soften the low pressure cut out tendancy.

5. Duke talked about having their remodel team come in-that is the people that do replace meters and regulators. TTC has worked with them before. Someone needs to be at the building with them while the boilers light off and show them what is happening. Also ask them to trend the gas pressure in the main, they have a device that does that.

Step 1- TTC will email Turton.

Step 2- TTC has already called Duke and will check back with them.

Step 3-Keep Boiler 1 in lead.

Step 4- A&M is coming out tomorrow.

10/18/11 – A&M fixed the software gliche on the boiler pumps on Friday October 14th.10/19/11 – The lighting board is not in yet and neither is the flame rod assembly but Ballinger will install when they come in. The gas piping has been changed and Boiler 2 went out overnight. Boiler 1 continues to run flawlessly. The errors are either #107 or #108 normally but we are going to send a hand written error log that Hamco has kept on Boiler 2.

10/19/11 – The lighting board is not in yet and neither is the flame rod assembly but Ballinger will install when they come in. The gas piping has been changed and Boiler 2 went out overnight. Boiler 1 continues to run flawlessly. The errors are either #107 or #108 normally but we are going to send a hand written error log that Hamco has kept on Boiler 2.

11/02/11 – Feldkamp had asked for final payment on this job. TTC e-mailed Pat Heeney to let him know that Hamilton County Facilities would not be willing to pay any more billings until Boiler #2 is fixed and acting properly. Currently it is going out every night, but seems to start back up in the mornings. Ballenger’s tech is out of town until Friday, but there was supposed to be an update on the flame rod replacement 2 weeks ago.

11/04/11 – TTC went to jobsite and witnessed the following: Flame rod sensor was changed out in both boilers. The new sensor is longer and thicker than the old one. BLR-2 started working properly as soon as the flame rod was changed out. The ignition start timing was changed to 2000. The low fire firing rate was changed to 1500 on both boilers.

TTC watched both boilers start up properly and it appeared that BLR-1 was back in LEAD and was getting ready to fire again at a setpoint at 150F.

11/07/11 – On Friday FLR-2 had 3 errors. One was error #108 and two were #109. The changes in the flame rod sensor and ignition timings were evidently not the answer. TTC requested a Tech visit and to call Mike Jackson at Hamilton Co. Facilities, he needs to know about the field visits since it is his building. No money will be released until this matter is resolved. Jim Turton with Ballenger Co. has advised Steve of the continuing problems at this site and will be in touch with Mike Jackson soon.

11/14/11 – Ballenger Company ‘s service department ordered a new boiler module for #2 and will install it as soon as it comes in. This should fix the issues. Ballenger will call the customer and schedule a service trip.

11/30/11 – The Boiler Module #2 was replaced on Monday 11/28/11, the combustion checked, and normal operation observed with no fault trips.

12/06/11 – Boiler #2 was out again. It went out on code 108-Flame Lost in Run. It went out yesterday (12/5) also on the same code.

12/09/11 – Steve Schill (service tech for Ballinger) will be at 250 WHT on 12/12 to run some test sequence operations. He has been in contact with Cleaver Brooks product engineers about the problems at this installation, and we will continue to work on a solution.

12/15/11 – TTC proposed a troubleshooting strategy that would allow tracking the boiler pressure at BLR-2 and BLR-1 and see what the differences are between the boilers and track those readings vs. failures vs incoming gas pressure from Duke. Either TEI does this survey themselves or they will hire a third party. If not problem presents itself TTC recommends the following changes at 250 WHT:

Ask to remove the pressure regulator inside the building and increase the pressure to the gas main to the building beyond the standard house pressure of 7-14″. The factory gas pressure regulators can be reinstalled and allow pressure to add one inline regulator on it too so we can drop its incoming pressure down to 14″. We could turn off permanently the natural gas feeding the labs or install a regulator on that line too but those may be out of service. This may be the best solution moving forward. If all agrees with the strategy, a face to face meeting should occur at 250 between Feldkamp, Ballenger, County and TEI.

12/20/11 – TTC spoke with Dave Messmer at Duke Energy and he says they can increase the gas pressure in the building but they would prefer to deliver 2 psig to the building not 1 psig. TTC asked Jim if the factory regulators from CB handle the 2 psig to 14″ wc turndown and if he sees any issue with it?

Assuming all that works, the County has two options at this point. Put in one 2 psig to 14″ regulator after the Duke regulator and make the whole building house pressure again or we can add regulators on each piece of equipment that doesn’t have them. Jim agrees that bringing 2 psig gas pressure into the building and then reducing it with a main step down regulator would be the “cleanest” approach. Instead of dropping the pressure to 14″ we, he suggest maybe consider a reduction of 9″ to 10″ WC in order to have a little safety factor against operating the boiler gas pressure regulators at what He believes is their maximum pressure limit. Ballenger’s service dept. needs to be notified 10 days before increasing the gas pressure.

12/22/11 – Here are the loads: We have two Cleaver Brooks Clearfire High Efficiency boilers, model number CFC-1000, Input 1000 mbh, Output 920 mbh. Only 1 of these boilers should ever run to handle the building load. The other boiler is 100% backup. Does it matter to Duke what the operation is of the boiler plant or do you size for peak load of all connected equipment? One water heater 250,000 btuh. All of these appliances are within 100 ft of the meter. Please let me know what the next step for increasing the pressure to 2 PSIG as you suggested is and what the schedule of that change might be. I have to have some contractors on site to add some regulators onto our equipment because they can only handle 14 psig incoming gas pressure so I have to coordinate that work too. Finally does Duke have a gas pressure trend log? I need to record the gas pressures before and after these changes.

12/29/11 – I have asked Duke to increase the pressure to the building to 2 psig. We will then have to install a new regulator on your house line that takes it from 2 psig down to 14″ but Jim Turton said he doesn’t want more than 10″wc which I struggle with because after we give him 10″ he will say in the dead of winter the boilers are going out on low gas pressure because Duke demand is so high. Anyways, it will probably be 12″. Their regulators can handle 1 psig (28″sc). I had to give all the loads in the building to Duke so they know what “new” regulator to bring out to the building. I have to confirm with Feldkamp that they will be purchasing and installing a new house regulator using the remaining allowance money. I think this will all take a little bit of time and with the holiday, I don’t think Duke is in a big hurry to get it done this week. What will be interesting to see is what happens when the new pressure to the BLR-2 unit is 10-12″. If it continues to g out an alarm then you will have a real case to make Feldkamp or Ballenger reimburse us all the cost or make it a no charge. Feldkamp has offered to swap the inline gas feeds between the boilers for troubleshooting – is that something you think is worth the effort if we are going with the 2spig to 12 psig regulator?

01/13/12 – TTC talked to Feldkamp and confirmed the new regular type and load. They are calling Duke to coordinate install in a week or two when it gets back up in the 40’s.

01/19/12 – The regulator and materials have been delivered to site and depending on the weather will be installed on Monday the 23 or Tuesday the 24. Start time will be around 9 or 10 to give the building time to warm up.

02/03/12 – Ballinger has been trying to get a factory tech in from Cleaver Brooks to resolve the issue. They hope to have this taken care of soon. Ballinger is copying a division head and V.P. on all correspondences.

02/21/12 – According to Jackson, the BLR-1 at 250 WHT made it through the weekend with no reported anomalies. Both boilers had the grounding point and wiring modified so they are both installed exactly the same way now.

03/16/12 – The County has accepted this project as completed. Terry has requested all closeout documents or an update on the status of the closeout process for a timeline for final payments from Pat at Feldkamp Enterprises.

03/28/12 – Terry C sent Pat at Feldkamp an e-mail saying the last payapp needs to be modified and the change order monies be paid. Terry requested the cost of change order for the extra meter at the building. There is $3,500 remaining allowance monies for this project. The previous change order was not valid so that needs to be torn up and Payapp #6 needs to be changed. Also Terry requested the final markups to create as-built drawings to submit to the County.

04/02/12 – The final invoice sent by Feldkamp is not payable through the County Auditor system, TEI will have to reject the entire pay app. Feldkamp needs to send a revised final payapp #6 that asks for the remaining balance including retainage which is $14,650.00. Send 3 originals to TTC notarized and signed for closeout.

Outstanding items that are required:

-AS-Built Markup – have they been submitted to show field changes?

-Final approved permit document from City

-Final payapp with correct billing amount

-Final certified payroll

-Three O&M manuals for Tony Matre

07/02/12 – TTC asked Pat Heeney for the status of the permit from the City? Are they going to accept our proposition? and that you have 2 weeks to turn in certified payroll to Bert Watts at HamCo before he turns this job into the State Wage and our Division. TTC will follow up with a phone call on status.

07/09/12 – Feldkamp talked with city inspector and looks like they will be ordering double wall vent pipe, not going to accept the way it is unless the manufacturer backs it up in writing which they won’t du to liabilities. Pretty sure there is no allowance monies left. Feldkamp will get measurements and get it on order.


250 WHT – Facade Assessment

THP Project Number: 11107.00

Contractor/HC Contract: NA

Contractors: NA

THP Documents: TBD

Contract Complete: NA


02/23/12 – SSRG completed test repair on west elevation.

04/25/12 – THP checking on status of leaks.

05/30/12 – THP checking on status of leaks with Mike Jackson.


250 WHT – General Health Alternate Study

Project Lead: GBBN

02/08/11 – GBBN met with Bert Watts and the staff from Public Health. The design development design and estimate were modified slightly to combine the Base and Alternate projects together into one project.

12/09/11 – GBBN sent the modified drawings and estimate based on the meeting held on 2/8/11

11/17/10 – Owner review of Design Development drawings and estimates. Owner required a modification of the SOSS to show an Alternate Design for the countertop computer modules.

11/18/10 – GBBN submitted a revised F021 and a new F023 SOSS to Bert Watts.

11/23/10 – GBBN sent final design development sketches and associated Opinions of Probable Costs to the owner.


250 WHT General Health Renovations

Project Lead: GBBN

07/28/10 – Attended Kick-off meeting with Bert Watts.

08/06/10 – SOSS sent to Bert Watts for approval.

09/20/10 – Meeting with Tim McCellan of Public Health to redefine the scope and change to two SOSS for alternate design options.

09/22/10 – Two SOSS sent to Bert Watts for approval.

11/17/10 – Owner review of Design Development drawings and estimates. Owner required a modification of the SOSS to show an Alternate Design for the countertop computer modules.

11/18/10 – GBBN submitted a revised F021 and a new F023 SOSS to Bert Watts.

11/23/10 – GBBN sent final design development sketches and associated Opinions of Probable Costs to the owner.

02/08/11 – GBBN met with Bert Watts and the staff from Public Health. The design development design and estimate were modified slightly to combine the Base and Alternate projects together into one project.

02/09/11 – GBBN sent the modified drawings and estimate based on the meeting held on 2/8/11.


250 WHT Chilled Water Pump Replacement

Project Lead: ThermalTech

03/14/13 – Background checks for A&M and Atkins & Stang are all approved. Two of three for Triton are approved.
03/18/13 – Triton is sending over project schedule for the job today. They are ready to start next week. The pumps ship on Friday. Brian wants to make sure they can drain down the chilled water and condenser water as needed for this work, TTC told them yes but they still have to coordinate with Jackson and Perkins.
03/25/13 – Triton sent TTC a cost sheet for replacing the existing 6″gate valve with a new Butterfly valve. TTC reviewed all work with Greg and Brian from Triton. They made a small piping change in the condenser water connections which TTC approved on site. The triple duty valve will stay in the system and be rebalanced to match the flow of the new CWP-1 being installed. The gate valve can be replaced for under $400 by Triton. Tony needs to give approval to allocate allowance money for this valve. Triton also noted that the pipe sizes on the Public Works documents are incorrect. The second floor piping is actually 4″ not 3″ and the first floor is actually 3″ not 2.5″. The flow and pressure drops are still fine. Triton is to mark up the As-Built drawings and they will be updated at the end of the job.

03/26/13 – Mike Jackson says all demo’d pumps/motors/rails, etc. needs to go. The only motor to stay will be the new motor (spare) that was purchased last year.

04/17/13 – The installed VFD and disconnect is acceptable for use on this project and meets the safety requirements and lock/tagout protocols of Hamilton County.

04/23/13 – A&M will be on site 4/24 to remove the existing DX-9100 expansion modules that presently control the chiller plant. They will then install a new BacNet controller to control the new chiller plant and tie it into the existing BacNet bus with the new VFD’s and the existing BacNet controller that was installed for the boiler project a couple of years ago. This will allow all of the chiller programming to be in one controller adjacent to the chiller. The intention is to have the chiller plant operational by the end of the day 4/24.

Triton sill try to have the balancer on site 4/24 so A&M can set up the controls with the minimum design flow of 110 GPM through the chiller.

04/24/13 – TTC and MJ review controls and mechanical work for pre punchlist work. Chiller is almost ready to be started.

04/26/13 – AHU-4 is low on flow-probably because of trapped air. CWP-1,2 are labeled backwards from dwgs and the head of the system are lower than designed thus increasing the flow readings on the most of the equipment. This is not the final balance report but is shows that we seem to have enough head and flow in the equipment to meet the design conditions. The final report will be included in the final O&M with the final punchlist for this building. Now waiting on controls update from Roger.

04/26/13 – A&M is still working on final checkout, so they did not have the DP sensors up and functional yet. Jim was going to let TTC know what they were reading so the setpoint could be determined. Without the system control, A&M was not able to place it into minimum flow condition (no control over the valves yet and the pump was in hand). Greg was aware as well, but A&M feels there is air in the lines that feed AHU-4 (east unit on the 2nd floor). It’s flow was much lower than the others regardless of where the pump was run. Greg is going to have maintenance look for air vents. Ultimately A&M still needs to verify AHU-4 flow after venting, verify final DP setting, and check flow through the bypass at AHU-1 with all coils closed. This work won’t take long, but needs to be done to ensure everything is ready before A&M schedules.

05/01/13 – The chiller is running and down to one override for now. TTC visited the site and walked the system with Will and talked to Roger on the phone about the programming. The chiller is up and running and with AHU-1,2,3 in automatic. However we need to keep AHU-4 in 85% open mode to meet the minimum flow of the chiller from field testing. From what TTC saw the dP across the chiller needs to be about 1.8 yesterday to keep it running. The mfg. data says a dP across the evaporator barrel only needs to be 0.9 for 116 gpm. TTC left the AHU-4 at 85% open and the dP at 1.8 when he left. This makes the VFD run about 64% speed which is pretty good at least there will be energy to save. On the flow issue there are a couple of items: The first issue is that for some reason there are multiple circuit setters that are not needed and they are installed in a bad location for our needs. The second issue is that we cannot easily replace the Hoffman valves at AHU-4 to bleed the air out of the system. Triton can replace the existing ones above the reception area for $988. TTC would prefer to have the County replace them instead and save the money for the water heater change order. The equipment is only $138 each. Is this a possibility? Triton to send a cost sheet over for review. There is a county meeting to discuss all options.

05/01/13 – The final report for the controls checkout and verification was completed and as follows:

AHU-4 has a 2 1/2″ circuit setter installed vs the 2″ installed at all other units. As the valves were already fully insulated when we arrived, I only cut access to the balancing ports and the position indicator. After trying to bleed air from the system, manipulating the other AHU control valves, and increasing the pump speed, Jim Biele and I were not able to get the coil to meet design. It was at the time I removed insulation to be sure we didn’t have some unseen problem, and found that the valve was in fact a 2 1/2″ valve. Using the proper performance data, we had adequate flow with the system at the original setpoint of 9 PSID.

The setpoint was determined as follows: pump at design flow, all AHU’s commanded to full cooling, all coils verified to be at design flow. The DP at that time was established as the setpoint. Upon completion of the venting and balancing on Monday, both DP sensors were reading the same value with the system at max flow. DP across the chiller at that time was approx 5.4 psid.

The final step was to verify minimum flow in the system with all valves closed. With AHU-1 in bypass (and the small unit in the basement that also has a 3 way control valve), AHU-2,3,4 closed, and the pump in auto 9 psid, we only measured about 55 gpm through the bypass of AHU-1. Jim increased the pump speed to see what would be required to achieve approx 110 gpm at the AHU and we were up to a DP of approx 25 psi before we met minimum flow (and the pump was at 85% at that point). I don’t see how we’re goint to meet the minimum flow requirement without another 3 way valve.

The bypass is piped such that there is a circuit setter (fully open presently) in the bypass line, and then a common return circuit setter downstream of the 3 way control valve. I had to use that common valve to set the flow through the coil. It is throttled to a setting of 14 (out of 100). Fully open, the coil would have too much flow but we would then have greater flow in bypass mode.

05/16/13 – The new bypass valve has been ordered. It’s a two week lead item. A&M should be installing in June and then TTC will commission it after it’s up and running. The plan is to restrict the bypass on AHU-1 at its circuit setter to a minimum and just use the new bypass as it is all we should need.

05/30/13 – CHW Bypass was approved, Triton waiting on change order approval for water heater. Final change order work should start in 2 weeks.

06/19/13 – ThermalTech meet with the Tony and Mike. They have concerns and all change order work is on hold until the bypass valve issue is resolved and working. ThermalTech contacted the contractors and advised them that more work needed to be done prior to installing the water heater.

06/24/13 – ThermalTech visited the site and collected valve positions, space temps, and differential pressure readings for the bypass work. All space temps appear normal.

08/22/13 – TTC met with Tony and Mike to present two options for modified control. Option 1 is a main 3-way valve and Option 2 is to retrofit the system with 3-way valves. We agreed to do Option #2.

08/23/13 – TTC approved 3-way valve installation with Triton. New valves are on order. TTC working with Triton and A&M for schedule and timing.

09/26/13 – Triton and A&M will install valves this week and commission system for operation to finish this project.

10/21/13 – 3-way valves installed. Water heater replacement alternate next. Then this project will be complete.

10/27/13 – Jackson reports all seems to be working very well.

12/13/13 – Startup domestic water heater. Denny & Triton on site for final startup. TTC provided some final installation instructions.

01/24/14 – Final paperwork being received by Triton for closeout. Signed change order came in for records.

01/30/14 – Need certified payroll and final closeout from all three contractors.

02/21/14 – Triton contact for closeout pay applications. TEI approved second to last payapp.

03/25/14 – Triton and TTC settled on May 1 warranty date. Final closeout paperwork enroute.

04/14/14 – Project complete final PW rates being sent to John Nester.


250 Energy Star 2014

Project Lead: TTE

10/09/14 – Terry and Bracken visited the building and took the air, people and light readings for the verification process.

10/23/14 – Final submission paperwork started but not completed yet. TTC is working on final submission for Ralph’s Signature.

11/28/14 – Building was approved for Energy Star with a score of 76 on Application # APP-20141119-0-1640490.


250 WHT Site Lighting

Project Lead: ThermalTech

02/26/12 – ThermalTech met with Will Perkins, Bert Watts and John Nester to review existing conditions and to define scope. ThermalTech developing SOSS.3/1/13 – SOSS issued to Bert Watts. ThermalTech waiting on approval and on PO.

03/01/13 – SOSS issued to Bert Watts. ThermalTech waiting on approval and on PO.

04/24/13 – SOSS approved and PO issued to ThermalTech. ThermalTech to schedule kick Off meeting with Bert Watts to establish schedule.

06/26/13 – Project bids have been received, and Debra Kuempel is apparent low bid. ThermalTech has called DK for pre-award conference, and will draft contracts as soon as conference is complete.

8/28/13 – DK has signed and returned contracts to Hamilton County. Waiting on NTP to begin work.

09/25/13 – DK has NTP – and will send Mike Barth of TE Submittals for approval. Tom Helmes of DK is matching existing equipment as specified.

10/30/13 – New pole has been received by DK and is scheduled to be installed this week.

01/24/14 – ThermalTech (Mike Barth) and DK met on site for final punchlist backcheck. DK submitted 100% pay application less retainage.

02/25/14 – DK submitted pricing for change order for replacing (6) lamps in parking lot, (1) wall pack lamp on building, and (3) pole light lamps at the coroners office. CO pending approval and signatures.

03/26/14 – DK has approved CO for replacing the lights.

04/24/14 – Change order sent in but not all the work was done so Mike and John need to coordinate the change order work. Ameresco lighting contractor jumped out and replaced lights.


250 WHT Penthouse Screen Wall Removal

Project Lead: THP

07/01/14 – THP issued review letter of penthouse conditions. County requested SOSS for 2014 structural services.

07/09/14 – THP issued SOSS to remove and repair penthouse structure.

07/17/14 – County approved SOSS.

07/22/14 – THP and County had a design coordination meeting.

07/31/14 – THP issued owner review documents.

08/04/14 – County provided review comments to project manual.

08/20/14 – THP and building tenants had meeting to discuss antennas located on the screen wall.

09/24/14 – THP issued bid document to County for review.

09/25/14 – THP issued bid documents to contractors for pricing.

09/26/14 – Bid Documents Issued.

09/30/14 – ThermalTech and THP met on site to discuss MEP options.

10/13/14 – THP issued addendum #1.

10/14/14 – Addendum No. 1 Issued.

10/21/14 – American Façade, Kramer, and O’Rourke submitted bids. American Façade was the apparent low bidder at $49044.00. THP qualifying bid.

10/28/14 – Bids received by American Façade Restoration, Kramer & Feldman, & O’Rourke. AFR is the apparent low bidder.

11/04/14 – THP reviewed and recommend AFR bid.

11/06/14 – THP issued draft AIA County / AFR contract.

01/07/14 – AIA Contract executed by prosecutor.

01/21/15 – Pre Construction Meeting completed.

01/22/15 – THP issued permit set to AFR and the County. County and THP reviewed and approved AFR bid.

01/28/15 – City of Cincinnati issued a building permit.

01/28/15 – All submittals returned to AFR.

02/03/15 – Preconstruction meeting completed.

02/18/18 – Review meeting for alternate antenna installations.

02/24/15 – THP provide options and pricing for alternate antenna installations. County is reviewing the options.

02/24/15 – All electrical work complete by Phasor Electrical Services with the exception of the reinstallation of the HAMM radio antennas. Antennas will be reinstalled once the screen walls are removed.

02/26/15 – Previous demolition work on hold due to weather. Work planned for 2/28/15, but waiting on final weather report.

03/14/15 – AFR removed the precast panels.

03/15/15 – AFR removed the steel and installed the antenna mast.

03/20/15 – AFR completed remaining work and THP published a punch list.

03/20/15 – City inspector issued a Final Inspection ticket.

04/28/15 – THP received final closeout paperwork form AFR and forwarded AFR closeout paper work to the County.

05/28/15 – THP to finish up CAD drawing and submit to the County.

06/24/15 – THP completed Archibus drawing files and AFR working on closeout project manual.

07/29/15 – Correction from 6/24, THP completed Archibus drawing files and AFR working on closeout project manual.


250 WHT Facade Repair

Project: F39

06/01/12 – THP submitted SOSS.

09/13/12 – County approves SOSS.

10/04/12 – THP County kick-off meeting complete. DD documents due early November and Final bid documents due mid-December with a budget. Bid period mid-February.

10/31/12 – THP working on DD documents.

01/15/13 – THP Issued DD documents for review.

02/08/13 – County returned project manual comments.

02/19/13 – THP and County had a design review meeting. Reviewed construction layout and phasing building needs. County requested additional window information for review.

05/21/13 – THP issued 90% Review documents to the County.

06/27/13 – Project update. Due to weather sensitive work, if project approved work will not occur until next spring. Possible bid project late 2012.

10/31/13 – Project on hold.

01/30/14 – Project on hold.

02/27/14 – Project on hold.

04/24/14 – Project on hold.

04/24/14 – County requested THP review the status of the penthouse precast panels.

05/23/14 – THP reviewed the cooling tower penthouse structure. Based on our observations the deterioration in the steel members has continued to increase. A letter will be issued with the results of the review along with our repair recommendations.

05/29/14 – Overall project on hold.

06/12/14 – THP completed a review of the penthouse cooling tower screen walls. The condition of the walls required stabilization emergency stabilization.

06/13/14 – THP facilitated a quick bid of stabilization of the panels and cooling tower beams. American Restoration was the only bidder able to respond quickly and there bid was within 10% of the estimated work.

06/14/14 – American Façade installed cables and shoring as directed by THP at the screen walls.

06/26/14 – THP anticipates publishing a letter of findings and recommendations by June 27th.

06/26/14 – Overall project on hold.

07/31/14 – Overall project on hold.

08/27/14 – Overall project on hold.

09/25/14 – Overall project on hold.

01/29/15 – Overall project on hold.

02/26/15 – Overall project on hold.

03/11/15 – THP assisted the County with the evaluation of falling concrete from the entry bridges.

03/12/15 – THP issued sketches to stabilize the bridge railings and allow use of the bridges.

03/16/15 – THP published an updated assessment of the entry bridges per the County’s request.

03/26/15 – Overall project on hold.

04/13/15 – County requested an SOSS to repair/replace the bridges. See F66.

04/30/15 – Project been put on hold.

12/09/15 – Projected construction 2017.


250 WHT Window Replacement

Project Lead: THP

06/01/12 – THP submitted SOSS.

09/13/12 – County approves SOSS.

10/04/12 – THP County kick-off meeting complete. DD documents due early November and Final bid documents due mid-December with a budget. Bid period mid-February.

10/31/12 – THP working on DD documents.

01/15/13 – THP Issued DD documents for review. Documents part of façade repair project as alternate.

09/25/13 – REFER to F39 for project status.

12/09/15 – Project on hold. Projected construction 2017


250 WHT Energy Star

Project Lead: Active

10/13/15 – TTC visited the site for energy star verification. Everything looked good.

12/08/15 – Energy Star Certification under consideration.


250 WHT Door Access Control

Project Lead: GBBN

01/06/16 – GBBN receives email, with scoping sketch attached, about project request. Facilities requests SOSS.

01/11/16 – GBBN issues two SOSS, one for each option, to Facilities Department.

04/18/16 – Facilities sends Estimating Project Requests showing this project as one of the listed projects.

06/30/16 – Submitted as a 2017 Project request.


250 WHT Bridge Replacement Special Inspection

Project Lead: THP – F79

06/06/16 – As part of the plans review the city required special inspections for both concrete testing and soil investigation.

06/14/16 – THP submitted an SOSS with Terracon completing the special inspection.

06/21/16 – County approved SOSS.

08/25/16 – Refer to F66 for updates.

01/26/17 – City requires extra paperwork from Terracon before they can final the permit on this job.

07/27/17 – City closed out permit for this project. All wrapped up. Ready to Archive.


250 WHT Bridge Replacement

Project Lead: THP – F66

04/13/15 – County requested SOSS to replace the bridges.

04/28/15 – THP issued SOSS. BW says the funding is approved for this project. PO forthcoming.

05/29/15 – County executed the SOSS.

06/17/15 – THP issued review documents to the County.

07/15/15 – THP met with City Fire Department plans examiner to confirm egress requirements.

07/22/15 – THP, Facilities department, and water department met to discuss project impact on users.

07/29/15 – THP, Facilities department, and public health department met to discuss project impact on users.

07/30/15 – THP anticipates issuing bid documents first week of August.

08/14/15 – THP issued bid documents.

08/24/15 – Pre-bid meeting completed.

08/26/15 – Project update: Bids are due 9/8/15.

09/02/15 – Addendum No. 1 Issued, changing the bid due date.

09/16/15 – Addendum No. 2 Issued modifying documents, issuing meeting minutes and answering contractor questions.

09/21/15 – Addendum No. 3 issued answering contractor question.

09/22/15 – Addendum No. 4 issued answering contractor question, issuing a new bid form, and changing the bid due date to 9/25/15.

9/25/15 – One bid submitted above the 10% allowable advertised budget.

10/01/15 – THP published bid finding letter and recommendation. Bid rejected by the County.

10/15/15 – THP published an updated budget based on current market conditions.

12/18/15 – County approved new budget and scope. THP authorized to finalized documents and bid the project.

01/11/16 – THP published updated review drawings and met with County to discuss documents.

01/20/16 – THP issued repair documents for bidding. ITB #008-16.

01/25/16 – THP issued Addendum No. 1, which changed the prebid meeting time to 11 AM.

01/28/16 – Prebid meeting scheduled at 11 AM.

02/12/16 – Bids were received from Smith Construction, Sunesis, Empire, ZSR, Prus, and SSRG. Smith Construction was the apparent low bidder. The submitted bid was within budget.

02/25/16 – THP review the bid with Smith Construction and contacted the submitted references. Based on bid amount, discussion with Smith Construction, and the reference comments; THP recommends Smith Construction.

02/25/16 – THP delivered recommendation letter and draft AIA contract to the County for review.

04/13/16 – County commissioners approved the project. THP notified Smith Construction project is moving forward. Smith Construction planned on starting the work in June.

04/26/16 – County approved the Contract and Issued a PO. THP forwarded both the Contract and PO to Smith Construction.

05/09/16 – THP provide sealed drawings to Smith Construction for permitting.

05/27/16 – Smith Construction working through background checks and fire training.

06/24/16 – Project is currently waiting permit

06/24/16 – Submittals are in.

07/07/16 – Permit approved.

07/20/16 – Preconstruction Meeting Completed.

07/28/16 – Smith Construction finalizing submittals. Start of site work anticipated for August 1st. Next site meeting August 8th.

08/03/16 – Preliminary meeting with building occupants completed to discuss fire watch plan.

08/08/16 – Fire Watch meeting completed between fire department, building department manager, HC Safety, Smith Construction, THP, and County Facilities.

08/15/16 – Construction began for the project. Demolition of Bridge #3 complete and new footing poured.

08/22/16 – Construction progress meeting 02 completed.

08/25/16 – Anticipate pouring bridge #3 foundation walls today.

08/25/16 – Bridge #3 foundation was poured.

09/06/16 – Bridge #3 upper landing and steps poured.

09/12/16 – Bridge #3 lower landing poured.

09/14/16 – Bridge #3 sidewalk adjacent to lower landing poured.

09/15/16 – Bridge #3 concrete test cylinder breaks completed and passed.

09/16/16 – Bridge #3 reopened with temporary rails.

09/19/16 – Demolition of Bridge #1/#2 started.

10/03/16 – Change Order 1 drafted and sent out

10/04/16 – Wall footings poured by Smith without approval of owner’s special inspector

10/06/16 – Change Order 2 for additional sidewalk drafted and sent out

10/11/16 – Terracon approved the excavated bearing condition

10/12/16 – Building struck by Smith dump truck

10/13/16 – Pay App 1 sent to All parties

10/24/16 – Bridge 1 and 2 poured.

11/21/16 – Bridge 1 and 2 structural slab poured.

12/02/16 – Bridge 1 and 2 lower slab on grade poured.

12/08/16 – Bridge 1 and 2 sidewalk slab on grade poured. 12/12/16 – Bridge 1 and 2 opened with temporary rails.

12/19/16 – THP and Smith completed measurements of unit cost allowance repairs.

12/22/16 – Bridge 4 and stairs below were poured.

12/29/16 – Bridge 4 opened with temporary rails.

01/24/17 – Started installation of railings for Bridges 1, 2, and 4.

02/21/17 – Smith still waiting to receive guardrails for Bridge #1 and #2. Tentatively plan to start work on Monday 2/27 and complete on 3/2.

03/23/17 – Project update: Alumina installed all permanent railings. Punch list items planned to be completed this week.

03/24/17 – THP completed site visit and published a punch list as part of the field report.

04/25/17 – Terracon visit the site to complete final special inspection work. Smith Construction not on site, so Terracon left. Terracon visit to be rescheduled.

05/23/17 – THP sent draft final Change Order 03 to Smith Construction.

05/25/17 – All field work complete, including punch list.

07/19/17 – THP sent request to Smith Construction to split payment of CO#3 at the request of the county. Follow-up sent 7/25. No response from Smith yet.

08/25/17 – THP received Smith Construction’s closeout paper work and final pay application.

08/31/17 – THP reviewing closeout paper work and final pay application.

09/05/17- Smith sent final warranty paperwork. THP reviewing.

09/07/17 – THP processed all paper work and sent to County.


250 WHT Recycling Fencing Location

Project Lead: GBBN – F079

03/23/17 – At Partnering Meeting Facilities explains project scope and requests SOSS.

03/24/17 – GBBN corresponds with Bayer Becker requesting proposal for partial survey and staking services.

04/07/17 – Bayer Becker sends proposal to GBBN.

04/10/17 – GBBN authors and delivers SOSS to Facilities.

07/17/17 – Bayer Becker has completed all field work.  The next step will be completing the survey drawing and staking the property line for the fence installation.  This should be completed by the end of the month.

07/26/17 – Bayer Becker sends GBBN a sign and sealed survey for use by the Owner.  Missing monuments and two sets foe nails have been installed on site for use in locating the fencing by Others.

10/23/17 – GBBN receives invoice from Bayer Becker for survey.


250 WHT Entry Leak Investigation

Project Lead: THP

05/12/17 – Hamilton County requested SOSS to investigate leaking around the building entries.

05/16/17 – THP visited the site and met with Mr. Perkins to review the leaking conditions to develop a proposal.

05/25/17 – THP developing SOSS.

06/08/17 – THP issued SOSS for review.

06/26/17 – County approved SOSS.

07/27/17 – THP plans to schedule work second week of August.

08/01/17 – THP reviewed the interior conditions of known water leaks and performed water testing on one entry door system and successfully got the entry to leak. Additional testing was performed on the windows.

08/22/17 – THP assessed all public entrances for signs of water infiltration.

09/01/17 – THP issued draft report to County for review

09/06/17 – THP issued final report to County.


250 – General Health Door Renovation

Project Lead: GBBN
07/28/17 – Hamilton County requested SOSS to investigate adding a new door and new Office Wall.
07/31/17 – GBBN visited the site and met with General Health to discuss scope of work. A hand drawn sketch was presented to GBBN. GBBN photographed and field verified the site. GBBN and TEI verified the proposed scope of work would require permitting.
08/03/17 – GBBN authored and issued SOSS for review.
08/21/17 – County requested a revised SOSS, based on a reduced scope to be compliant with no permitting required. A simple opening through the existing gypsum board partition is requested. TEI confirmed no added load to electrical service means no IBI permit required.
08/23/17 – Facilities clarifies the final scope of the project to GBBN and TEI.
08/25/17 – GBBN authors and delivers revised SOSS to Facilities.
09/18/17 – GBBN clarifies scope with Facilities to use hollow metal door frame in new opening for durability. Double studs and header will be used in expanded opening rough framing for frame installation and anchorage.
09/22/17 – GBBN confirms to Facilities the drawings are completed to the point of final field verification and owner review of drawings.
10/02/17 – Facilities requests final drawings from GBBN for issuance to Trades. GBBN sends drawings with inquiry about final Owner review meeting.